OpenStudy (anonymous):

Write the equation in slope-intercept form of the line that has a slope of -3 and contains the point (4, -5).

OpenStudy (anonymous):

please help me with problem A door to door salesman averages one sale for each 6 contacts. What linear equation fits this situation?

OpenStudy (anonymous):

please and thank you

OpenStudy (dragons1415):

so you dont need the first one?

OpenStudy (anonymous):

no i dont im srry need answer to new one now

OpenStudy (anonymous):

can u please help

OpenStudy (anonymous):

answer choices are x=y+6, y=(1/6)x, y=x+6, x=6y

OpenStudy (dragons1415):

I don't know, sorry :C

OpenStudy (anonymous):

its ok

OpenStudy (dragons1415):

but it's either y=(1/6)x or x=6y

OpenStudy (anonymous):

it was y=(1/6)x thanks

OpenStudy (dragons1415):

yeah, no prob